OpenStudy (anonymous):

Will give medal to best answer and fan 2 gal/hour=____floz/min

OpenStudy (anonymous):

@JayJayTheWereWolfGal1

OpenStudy (mertsj):

\[\frac{2gal}{1 hr} \times\frac{1hr}{60\min} \times \frac{128 fl oz}{1 gal}=\]

OpenStudy (mertsj):

Grind it out

OpenStudy (anonymous):

whao...

OpenStudy (anonymous):

so 1 gal and 1128 oz over 60 min

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

128*128/60min?

OpenStudy (mertsj):

cancel the gal

OpenStudy (mertsj):

cancel the hrs

OpenStudy (anonymous):

i got 128*128/60min @Mertsj

OpenStudy (mertsj):

now you have: \[\frac{2 \times 1 \times 128floz}{60 \min}\]

OpenStudy (mertsj):

Now cancel the 2 into the 60

OpenStudy (anonymous):

128 fl oz/ 30?

OpenStudy (mertsj):

\[\frac{128 fl oz}{30 \min}\]

OpenStudy (mertsj):

=4.3 fl oz per minute

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

thanks